Show that the function f defined by f(x) =1//x^2 is uniformly continuous on [5 ,∞].

1
Expert's answer
2022-05-25T09:13:00-0400

ANSWER

Let "\epsilon >0." By the definition, we need to show there exists "\delta >0" , such that


"a\geq 5 , b\geq 5" and "|a-b|<\delta" imply "|f(a) - f(b)|< \epsilon" (1)

The function "f(x)=\frac{1}{x^{2}}" is differentiable on the set "[5,\infty)" . "f'(x)=\frac {-2} { x^{3}}" and "|f'(x)|\leq \frac{2}{5^{3}}<1" for all "x\in [5,\infty)." Let "b>a\geq 5" , according to the mean value theorem, there exists point "c\in (a,b)" such that "f(b)-f(a) =f'(c)\cdot (b-a)". So, for all "a,b\in [5,\infty)" we have

"|f(b)-f(a)|\leq |b-a|" .

Hence, if "\epsilon >0" and "\delta=\epsilon" , then

"|a-b|<\delta =\epsilon \Rightarrow |f(b)-f(a)|\leq |b-a|<\delta=\epsilon"

Those, (1) is satisfied, which means that the function is uniformly continuous.